# Characterizing Multimodal Long-form Summarization: A Case Study on Financial Reports

### v1

@@ 7 Conclusion
A.1 Cohere
Among all the 1000 summaries, 46.7% only generate “I’m sorry, but I am unable to complete your request because I am unable to identify the relevant information to complete the task.” This suggests that Cohere fails to deal with tasks in which the input length is almost equal to the context limit. However, to be noticed, Cohere is quite good at reading tables in its completed summary, where 30.69% of the numbers are extracted only from tables
A.2 GPT-3.5
@@ Appendix E Details of Similarity Score Calculating Algorithm
0.6-0.7
summary sentence: Boat segment sales declined due to production disruptions from covid-19, but rebounded in the second half of 2020.
0.67
@@
summary sentence: - Free cash flow was \$629 million in 2020, enabling debt paydown, share repurchases and dividends.
0.65
@@
0.7-0.8
summary sentence: - Total operating expenses increased 2.0% to \$42.7 billion, driven by higher salaries, maintenance, and regional capacity costs.
0.74
@@
summary sentence: - At December 31, 2016, nhi had \$2.6 billion invested in 205 facilities across 32 states.
0.73
@@
0.8-0.9
summary sentence: - American reached a confidential settlement with Boeing regarding financial damages from the 737 max grounding.
0.85
@@
summary sentence: - Cash from operations dropped to \$209 million in 2019 from \$389 million in 2018 due to lower earnings.
0.82
@@
0.9-1.0
summary sentence: Passenger revenue was \$42.0 billion, up 3.3% due to higher passenger demand.
0.92
@@
summary sentence: - Company restaurant revenues come from food and beverage sales at company-operated restaurants.
0.94
@@
>1.0
summary sentence: Net earnings were \$374.7 million in 2020 versus \$30.4 million in 2019.
1.13
@@
summary sentence: Franchise royalty and other franchise revenues come from royalty income and initial and renewal franchise fees.
1.38

### v3

@@ Acknowledgements
A.1 Cohere
Among all the 1000 summaries, 46.7% only generate “I’m sorry, but I am unable to complete your request because I am unable to identify the relevant information to complete the task.” This suggests that Cohere fails to deal with tasks in which the input length is almost equal to the context limit. However, to be noticed, Cohere is quite good at reading tables in its completed summary, where 30.69% of the numbers are extracted only from tables.
A.2 GPT-3.5
@@ Appendix D Prompts
Chain-of-Thought (CoT) prompt
{spverbatim}
Summarize the following report.
MD&A: …
Let’s generate the summary step by step.
1. Read through the entire MDA report carefully to understand the context. 2. Identify and extract the key topics and insights discussed in the report. 3. Pay attention to any tables presenting numeric data, such as income statements, balance sheets, or cash flow statements. 4. When including numbers in the summary, ensure they are: a) Explicitly stated values from the original report (do not fabricate numbers). b) Stemmed from step-by-step verified calculations. c) Correctly rounded. d) Appropriately represented with clear context from the original source. 5. Synthesize the extracted information and numbers into a concise summary that flows logically.
Summary:
@@ Appendix E Details of Similarity Score Calculating Algorithm
0.6-0.7 summary sentence: Boat segment sales declined due to production disruptions from covid-19, but rebounded in the second half of 2020. 0.67
@@
summary sentence: - Free cash flow was \$629 million in 2020, enabling debt paydown, share repurchases and dividends. 0.65
@@
0.7-0.8 summary sentence: - Total operating expenses increased 2.0% to \$42.7 billion, driven by higher salaries, maintenance, and regional capacity costs. 0.74
@@
summary sentence: - At December 31, 2016, nhi had \$2.6 billion invested in 205 facilities across 32 states. 0.73
@@
0.8-0.9 summary sentence: - American reached a confidential settlement with Boeing regarding financial damages from the 737 max grounding. 0.85
@@
summary sentence: - Cash from operations dropped to \$209 million in 2019 from \$389 million in 2018 due to lower earnings. 0.82
@@
0.9-1.0 summary sentence: Passenger revenue was \$42.0 billion, up 3.3% due to higher passenger demand. 0.92
@@
summary sentence: - Company restaurant revenues come from food and beverage sales at company-operated restaurants. 0.94
@@
>1.0 summary sentence: Net earnings were \$374.7 million in 2020 versus \$30.4 million in 2019. 1.13
@@
summary sentence: Franchise royalty and other franchise revenues come from royalty income and initial and renewal franchise fees. 1.38